Seafoam:

A triangle has sides that measure 7 units, 6 units, and 9.22 units. What is the area of a circle with a circumference that equals the perimeter of the triangle? Use 3.14 for π, and round your answer to the nearest whole number. 22 units2 157 units2 39 units2 70 units2

Seafoam:

39

Seafoam:

@AZ

bigblood:

Yeah Can I be in this also

Seafoam:

Haha sure, I'm just running my answers through here to check

bigblood:

Ok

AZ:

okay so circumference = 7 + 6 + 9.22 22.22 = 2pi * r r = 3.53 area = pi * r^2

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Seafoam 39 \(\color{#0cbb34}{\text{End of Quote}}\) and so yeah, I came to the same answer